# Non-linear conductances of Galton–Watson trees and application to the (near) critical random cluster model

### v1

@@ 2.1 The random cluster model on a graph
where $o(\omega)=\sum_{e\in E}\omega_{e}$ is the number of open edges, $f(\omega)=\sum_{e\in E}1-\omega_{e}$ is the number of closed edges and $k(\omega)$ is the number of clusters in $\omega$, i.e. the number of connected components of the subgraph of $G$ induced by $\omega$. Here, $Z_{\mathtt{p},\mathtt{q}}^{G}$ is the partition function of the model, that is the constant that normalizes $\mathbb{P}_{\mathtt{p},\mathtt{q}}$ to a probability. Let us mention that when $\mathtt{q}\geqslant 1$, the model enjoys some monotonicity and correlation inequalities, see e.g [21, Ch. 3]; these are essential tools for many of the techniques developed for the study of the RCM, which is therefore often considered only for $\mathtt{q}\geqslant 1$.
Relation with percolation and Ising/Potts models.
Notice that for $\mathtt{q}=1$, one clearly recovers the percolation model with parameter $\mathtt{p}$. On the other hand, for $\mathtt{q}\in\{2,3,4,\ldots\}$ the RCM can be coupled to the (ferromagnetic) $\mathtt{q}$-states Potts model with inverse temperature $\beta$ verifying $e^{-\beta}=1-\mathtt{p}$, see [21, §1.4], defined by the Gibbs measure on configurations $\sigma\in\{1,2,\ldots,\mathtt{q}\}^{V}$ by
@@
Additionally, let us stress that in the coupling between the RCM and the $\mathtt{q}$-state Potts model, one has a direct relation between the two-point correlation functions for the Potts model and connection probabilities in the RCM, see [21, Thm. 1.16]:
@@
Boundary conditions.
Analogously to the Ising and Potts models, one may introduce a boundary condition when considering the random cluster model. Here, we will only focus on the wired boundary condition on the graph, which consists in contracting all vertices of the boundary $\partial G$ into one single vertex denoted $\{\partial G\}$: we denote by $\bar{G}$ the resulting graph and ¯$\bar{\mathbb{P}}_{\mathtt{p},\mathtt{q}}^{\bar{G}}$ the resulting RCM. In terms of the Potts model, this corresponds to putting all spins equal on the boundary; for the Ising model this is the model with plus (or minus) boundary condition.
@@ 2.2 The random cluster model on trees and its associated recursion
is one of the key quantity of interest; note that it depends on the realization of the tree $T$ and we will treat it for a quenched realization of $T$. In particular, in the case where $\mathtt{q}=2$, one has that $\pi_{n}$ is equal to the magnetization of the root in the Ising model on $T_{n}$ with plus boundary condition on the leaves $\partial T_{n}$; a similar interpretation holds for $\pi_{n}$ in the $\mathtt{q}$-Potts model if $\mathtt{q}\in\{2,3,\ldots\}$, see (2.2). In particular, one is interested to know whether, depending on the parameters $\mathtt{p},\mathtt{q}$, the connection probability $\pi_{n}$ vanishes or not as $n\to\infty$, and if it does, at which rate.
Let us mention that the RCM on $d$-ary (or Cayley) trees has been studied in several articles, see e.g. [21, §10.9-10] and references therein. In a nutshell, the results from [22] ([21] only treats the case $d=2$) state that for every fixed $\mathtt{q}>0$ a phase transition occurs at some $\mathtt{p}_{c}(q)\in(0,1)$ given by different formulas according to whether $\mathtt{q}\leqslant 2$ or $\mathtt{q}>2$: if $\mathtt{q}\in(0,2]$, then $\mathtt{p}_{c}(\mathtt{q})=\frac{\mathtt{q}}{d+\mathtt{q}-1}$ and the phase transition is continuous; if $\mathtt{q}>2$, then $\mathtt{p}_{q}(\mathtt{q})$ is the unique value of $\mathtt{p}$ such that the polynomial $Q_{\mathtt{p},\mathtt{q}}(x):=(\mathtt{q}-1)x^{d+1}-(\mathtt{q}-1+\frac{\mathtt{p}}{1-\mathtt{p}})x^{d}+\frac{1}{1-\mathtt{p}}x-1$ has a double root in $(0,1)$ (note that for $d=2$, this gives $\mathtt{p}_{c}(q)=\frac{2\sqrt{\mathtt{q}-1}}{1+2\sqrt{q-1}}$), and the phase transition is discontinuous.
The random cluster model seems to have been considered on $d$-regular (random) graphs, see e.g. [4, 5, 23], which are locally $d$-ary trees. However, it does not seem to have been studied on more random graphs with no fixed degree, in particular, to the best of our knowledge, it has not been studied on random trees (for instance Galton–Watson trees). Our goal is to show that the connection probabilities verify some recursive relation on trees and apply our results of Section 1.3 to obtain information on the critical (or near and sub-critical) random cluster model.
The RCM iteration on trees.
@@
$\varphi_{\mathtt{q}}(\pi_{n}(u))=\prod_{y\leftarrow u}\varphi_{\mathtt{q}}\big{(}\gamma\,\pi_{n}(v)\big{)}\quad\text{ where }\ \varphi_{q}(x):=\frac{1-x}{1+(\mathtt{q}-1)x}$ (2.3)
and $\gamma=\gamma_{\mathtt{p},\mathtt{q}}:=\varphi_{\mathtt{q}}(1-\mathtt{p})$. Setting $B_{n}(u):=-\log\varphi_{\mathtt{q}}(\pi_{n}(u))$, we have $B_{n}(u)=+\infty$ if $u\in\partial T_{n}$ and, if $u\notin\partial T_{n}$ then we have the recursion: letting $\beta$ be such that $1-\mathtt{p}=e^{-\beta}$,
$B_{n}(u)=\sum_{v\leftarrow u}\psi_{q}^{-1}\big{(}\psi_{\mathtt{q}}(\beta)\psi_{\mathtt{q}}(B_{n}(v))\big{)}\,,\quad\text{ where }\ \psi_{\mathtt{q}}(x):=\frac{e^{x}-1}{e^{x}+\mathtt{q}-1}=\varphi_{\mathtt{q}}(e^{-x})\,.$ (2.4)
Let us stress that when $\mathtt{q}=2$, we have that $\psi_{q}(x)=\tanh(x)$, so one recovers the well-known Lyons’ iteration for the Ising model, see [25] (see also [13, Lem. 2.3] for the specific form $\psi_{q}^{-1}\big{(}\psi_{\mathtt{q}}(\beta)\psi_{\mathtt{q}}(x)\big{)}=\tanh^{-1}(\tanh(\beta)\tanh(x))$ we have here).
@@
Let us now stress that the function $g_{\beta}$ is concave on $\mathbb{R}_{+}$ if and only if $\mathtt{q}\in(0,2]$; we will therefore focus on the case $\mathtt{q}\in(0,2]$. As a clue for this fact, notice that $g_{\beta}^{\prime\prime}(0)=\frac{\mathtt{q}-2}{\mathtt{q}}(1-\psi_{\mathtt{q}}(\beta))$, which is negative if $\mathtt{q}<2$, equal to zero if $\mathtt{q}=2$ and positive if $\mathtt{q}>2$.
@@
[…]
so we indeed have (1.5) with $q=2$ (and $s=1$) if $\mathtt{q}\in(0,2)$ and with $q=3$ (and $s=3$) if $\mathtt{q}=2$; this was noticed in [33] in the case of the Ising model (with plus boundary condition, which correspond to our wired boundary condition for the RCM).
@@ 2.3.1 The critical point and critical behavior at (and below) criticality
For any fixed $\beta$ (recall $e^{-\beta}=1-\mathtt{p}$), we have that (2.5) is exactly the recursion (1.8) with $R_{n}\equiv R=\psi_{\mathtt{q}}(\beta)$, and (1.5) holds with $q=2$ ($s=1$) if $\mathtt{q}\in(0,2)$ and with $q=3$ ($s=2$) if $\mathtt{q}=2$. First of all, Theorems 1.6-1.10 allows us to identify the critical point.
@@
$\pi_{n}(\beta)=\pi_{n}^{\mathtt{q}}(\beta,T):=\bar{\mathbb{P}}_{\mathtt{p}_{n},\mathtt{q}}^{\bar{T}_{n}}(\rho\leftrightarrow\partial T_{n})\,,\qquad\pi(\beta)=\lim_{n\to\infty}\pi_{n}(\beta)\,,$ (2.7)
@@
Recall that $B_{n}=-\log\varphi_{\mathtt{q}}(\pi_{n})$, so in particular $\pi_{n}$ goes to 00 if and only if $B_{n}$ goes to 00.
@@
Theorem 2.4 therefore identifies the critical point $\mathtt{p}_{c}=1-e^{-\beta_{c}}$ for the random cluster model with $\mathtt{q}\in(0,2]$ on a (random) Galton–Watson tree. The case $\mathtt{q}>2$ is understood in the case of a $d$-ary tree but remains somehow mysterious on a random tree and let us briefly comment on it. In fact, when $\mathtt{q}>2$ the phase transition on a $d$-ary tree happens at a certain $\mathtt{p}_{c}(\mathtt{q})$ characterized by being the only (double) root in $(0,1)$ of a certain polynomial $Q_{\mathtt{p},\mathtt{q}}^{(d)}(x)$ (see 2.2). In terms of the recursion (2.5), this correspond to finding $\mathtt{p}$ such that $dg_{\beta,\mathtt{q}}$ (see (2.2)) intersects the identity function exactly once. For the RCM on a Galton–Watson tree, if $\mathtt{q}>2$, the random recursion (2.5) is still valid but the function $g_{\beta}$ is then not concave anymore, and it is not clear whether a characterization of $p_{c}(\mathtt{q})$ in terms of a certain (averaged?) polynomial remains valid.
@@
The following result is a direct consequence of Theorem 1.8-Proposition 1.9 in the case where $Z\sim\mu$ admits a finite moment of order $q$ and of Theorem 1.10 in the case where $\mathbf{E}[Z^{q}]=+\infty$. Notice that $-\log\varphi_{\mathtt{q}}(x)\sim\mathtt{q}x$ as $x\downarrow 0$, so $\pi_{n}\sim\mathtt{q}^{-1}B_{n}$ when $B_{n}$ goes to 00. Additionally, note that from (2.6) we get that $\kappa_{g}:=\lim_{x\downarrow 0}\frac{1}{x^{2}}(x-g_{\beta}(x))=\frac{\mathtt{q}-2}{\mathtt{q}}(1-\psi_{\mathtt{q}}(\beta))$ if $\mathtt{q}\in(0,2)$ and $\kappa_{g}:=\lim_{x\downarrow 0}\frac{1}{x^{3}}(x-g_{\beta}(x))=\frac{1}{3}(1-\psi_{\mathtt{q}}(\beta)^{2})$ if $\mathtt{q}=2$.
@@
Let $\beta=\beta_{c}$, with $\beta_{c}$ defined by the relation $m\psi_{\mathtt{q}}(\beta_{c})=1$. We then have the following asymptotic results for $\pi_{n}$. Letting $q=2$ ($s=1$) in the case $\mathtt{q}\in(0,2)$ and $q=3$ ($s=2$) in the case $\mathtt{q}=2$, we have the following:
  • if $\mathbf{E}[Z^{q}]<+\infty$, we have, almost surely and in $L^{q}$,
    $\lim_{n\to\infty}n^{\frac{1}{q-1}}\pi_{n}=\alpha_{\mathtt{q}}W\qquad\text{ with }\ \alpha_{\mathtt{q}}:=\begin{cases}\frac{\mathtt{q}}{\mathtt{q}-2}\frac{m}{m-1}\mathbf{E}[W^{2}]^{-1}&\text{ if }\mathtt{q}\in(0,2)\,,\\[3.0pt] \frac{\sqrt{3}}{\sqrt{2}}\frac{m}{\sqrt{m^{2}-1}}\mathbf{E}[W^{3}]^{-1/2}&\text{ if }\mathtt{q}=2\,.\end{cases}$
  • if $c_{1}L(x)x^{q-\alpha}\leqslant\mathbf{E}[(Z\wedge x)^{q}]\leqslant c_{2}L(x)x^{q-\alpha}$ for some $\alpha\in(1,q]$, some slowly varying function $L(\cdot)$ and constants $c_{1},c_{2}>0$, then
@@
In fact, Section 1.3 gives much more precise results, that we describe informally as follows: letting $q=2$ if $\mathtt{q}\in(0,2)$ and $q=3$ if $\mathtt{q}=2$, we have
  • if $\mathbf{E}[Z^{q}]<+\infty$, we have $\lim_{n\to\infty}(m\psi_{\mathtt{q}}(\beta))^{-n}\pi_{n}=\alpha_{\mathtt{q}}(\beta)W$ for some constant $\alpha_{\mathtt{q}}(\beta)$;
  • if $\mathbf{E}[(Z\wedge x)^{q}]=n^{c+o(1)}$ for some constant $c$, then $C\geqslant(m\psi_{\mathtt{q}}(\beta))^{-n}\pi_{n}\geqslant n^{-c^{\prime}}$ as $n\to\infty$.

### v2

@@ 2.1 The random cluster model on a graph
where $o(\omega)=\sum_{e\in E}\omega_{e}$ is the number of open edges, $f(\omega)=\sum_{e\in E}1-\omega_{e}$ is the number of closed edges and $k(\omega)$ is the number of clusters in $\omega$, i.e. the number of connected components of the subgraph of $G$ induced by $\omega$. Here, $Z_{\mathtt{p},\mathtt{q}}^{G}$ is the partition function of the model, that is the constant that normalizes $\mathbb{P}_{\mathtt{p},\mathtt{q}}$ to a probability. Let us mention that when $\mathtt{q}\geqslant 1$, the model enjoys some monotonicity and correlation inequalities, see e.g [23, Ch. 3]; these are essential tools for many of the techniques developed for the study of the RCM, which is therefore often considered only for $\mathtt{q}\geqslant 1$.
Relation with percolation and Ising/Potts models
Notice that for $\mathtt{q}=1$, one clearly recovers the percolation model with parameter $\mathtt{p}$. On the other hand, for $\mathtt{q}\in\{2,3,4,\ldots\}$ the RCM can be coupled to the (ferromagnetic) $\mathtt{q}$-states Potts model with inverse temperature $\beta$ verifying $e^{-\beta}=1-\mathtt{p}$, see [23, §1.4], defined by the Gibbs measure on configurations $\sigma\in\{1,2,\ldots,\mathtt{q}\}^{V}$ by
@@
Additionally, let us stress that in the coupling between the RCM and the $\mathtt{q}$-state Potts model, one has a direct relation between the two-point correlation functions for the Potts model and connection probabilities in the RCM, see [23, Thm. 1.16]:
@@
Boundary conditions
Analogously to the Ising and Potts models, one may introduce a boundary condition when considering the random cluster model. Here, we will only focus on the wired boundary condition on the graph, which consists in contracting all vertices of the boundary $\partial G$ into one single vertex denoted $\{\partial G\}$: we denote by $\bar{G}$ the resulting graph and $\bar{\mathbb{P}}_{\mathtt{p},\mathtt{q}}^{\bar{G}}$ the resulting RCM. In terms of the Potts model, this corresponds to putting all spins equal on the boundary; for the Ising model this is the model with plus (or minus) boundary condition.
@@ 2.2 The random cluster model on trees and its associated recursion
is one of the key quantities of interest; note that it depends on the realization of the tree $T$ and we will consider it under a quenched realization of $T$. In particular, in the case where $\mathtt{q}=2$, one has that $\pi_{n}$ is equal to the magnetization of the root in the Ising model on $T_{n}$ with plus boundary condition on the leaves $\partial T_{n}$; a similar interpretation holds for $\pi_{n}$ in the $\mathtt{q}$-Potts model if $\mathtt{q}\in\{2,3,\ldots\}$, see (2.2). In particular, one is interested in knowing whether, depending on the parameters $\mathtt{p},\mathtt{q}$, the connection probability $\pi_{n}$ vanishes or not as $n\to\infty$, and if so, at what rate.
Let us mention that the RCM on $d$-ary (or Cayley) trees has been studied in several articles, see e.g. [23, §10.9-10] and references therein. In a nutshell, the results from [24] ([23] only treats the case $d=2$) state that for every fixed $\mathtt{q}>0$ a phase transition occurs at some $\mathtt{p}_{c}(q)\in(0,1)$ given by different formulas according to whether $\mathtt{q}\leqslant 2$ or $\mathtt{q}>2$: if $\mathtt{q}\in(0,2]$, then $\mathtt{p}_{c}(\mathtt{q})=\frac{\mathtt{q}}{d+\mathtt{q}-1}$ and the phase transition is continuous; if $\mathtt{q}>2$, then $\mathtt{p}_{q}(\mathtt{q})$ is the unique value of $\mathtt{p}$ such that the polynomial $Q_{\mathtt{p},\mathtt{q}}(x):=(\mathtt{q}-1)x^{d+1}-(\mathtt{q}-1+\frac{\mathtt{p}}{1-\mathtt{p}})x^{d}+\frac{1}{1-\mathtt{p}}x-1$ has a double root in $(0,1)$ (note that for $d=2$, this gives $\mathtt{p}_{c}(\mathtt{q})=\frac{2\sqrt{\mathtt{q}-1}}{1+2\sqrt{\mathtt{q}-1}}$), and the phase transition is discontinuous.
The random cluster model seems to have been considered on $d$-regular (random) graphs, see e.g. [4, 5, 25], which are locally $d$-ary trees. However, it does not appear to have been investigated on more general random graphs without fixed degree. In particular, to the best of our knowledge, it has not been studied on random trees, such as Galton–Watson trees. Our goal is to show that the connection probabilities satisfy a recursive relation on trees and to apply our results of Section 1.3 to obtain information on the critical (or near and sub-critical) random cluster model.
2.2.1 The RCM iteration on trees
@@
$\varphi_{\mathtt{q}}(\pi_{n}(u))=\prod_{y\leftarrow u}\varphi_{\mathtt{q}}\big(\gamma\,\pi_{n}(v)\big)\quad\text{ where }\ \varphi_{\mathtt{q}}(x):=\frac{1-x}{1+(\mathtt{q}-1)x}$ (2.3)
and $\gamma=\gamma_{\mathtt{p},\mathtt{q}}:=\varphi_{\mathtt{q}}(1-\mathtt{p})$. Setting $B_{n}(u):=-\log\varphi_{\mathtt{q}}(\pi_{n}(u))$, we have $B_{n}(u)=+\infty$ if $u\in\partial T_{n}$ and, if $u\notin\partial T_{n}$, the following recursion holds: letting $\beta$ be such that $1-\mathtt{p}=e^{-\beta}$,
$B_{n}(u)=\sum_{v\leftarrow u}\psi_{\mathtt{q}}^{-1}\big(\psi_{\mathtt{q}}(\beta)\psi_{\mathtt{q}}(B_{n}(v))\big)\,,\quad\text{ where }\ \psi_{\mathtt{q}}(x):=\frac{e^{x}-1}{e^{x}+\mathtt{q}-1}=\varphi_{\mathtt{q}}(e^{-x})\,.$ (2.4)
Let us stress that when $\mathtt{q}=2$, we have that $\psi_{\mathtt{q}}(x)=\tanh(x)$, so one recovers the well-known Lyons’ iteration for the Ising model, see [28] (see also [14, Lem. 2.3] for the specific form $\psi_{\mathtt{q}}^{-1}\big(\psi_{\mathtt{q}}(\beta)\psi_{\mathtt{q}}(x)\big)=\tanh^{-1}(\tanh(\beta)\tanh(x))$ we have here).
@@ 2.2.2 Properties of the recursion
Let us now stress that the function $g_{\beta}$ is concave on $\mathbb{R}_{+}$ if and only if $\mathtt{q}\in(0,2]$; we will therefore restrict our attention to the case $\mathtt{q}\in(0,2]$. As a clue for this fact, notice that $g_{\beta}^{\prime\prime}(0)=\frac{\mathtt{q}-2}{\mathtt{q}}(1-\psi_{\mathtt{q}}(\beta))$, which is negative if $\mathtt{q}<2$, equal to zero if $\mathtt{q}=2$ and positive if $\mathtt{q}>2$.
@@
 $\begin{split}\text{ if }\mathtt{q}\in(0,2)\,,&\hskip 17.00024ptg_{\beta}(x)=x-\frac{2-\mathtt{q}}{2\mathtt{q}}(1-\psi_{\mathtt{q}}(\beta))x^{2}(1+o(1))\\ \text{ if }\mathtt{q}=2\,,&\hskip 17.00024ptg_{\beta}(x)=x-\frac{1}{12}(1-\psi_{\mathtt{q}}(\beta)^{2})x^{3}(1+o(1))\end{split}$ (2.6)
so we indeed have (1.9) with $s=1$ if $\mathtt{q}\in(0,2)$ and with $s=2$ if $\mathtt{q}=2$; this was noticed in [36] in the case of the Ising model (with plus boundary condition, which correspond to our wired boundary condition for the RCM).
@@ 2.3.1 The critical point and critical behavior at (and below) criticality
For any fixed $\beta$ (recall $e^{-\beta}=1-\mathtt{p}$), we have that (2.5) is exactly the recursion (1.8) with $R_{n}\equiv R=\psi_{\mathtt{q}}(\beta)$, and (1.9) holds with $s=1$ if $\mathtt{q}\in(0,2)$ and with $s=2$ if $\mathtt{q}=2$. First of all, Theorems 1.6-1.10 allows us to identify the critical point.
@@
$\pi_{n}(\beta)=\pi_{n}^{\mathtt{q}}(\beta,T):=\bar{\mathbb{P}}_{\mathtt{p},\mathtt{q}}^{\bar{T}_{n}}(\rho\leftrightarrow\partial T_{n})\,,\qquad\pi(\beta)=\lim_{n\to\infty}\pi_{n}(\beta)\,$, (2.7)
@@
Recall that $B_{n}=-\log\varphi_{\mathtt{q}}(\pi_{n})$, so in particular $\pi_{n}$ goes to 0 if and only if $B_{n}$ goes to 0.
@@
Theorem 2.4 therefore identifies the critical point $\mathtt{p}_{c}=1-e^{-\beta_{c}}$ for the random cluster model with $\mathtt{q}\in(0,2]$ on a (random) Galton–Watson tree. The case $\mathtt{q}>2$ is understood in the case of a $d$-ary tree but remains somehow mysterious on a random tree and let us briefly comment on it. In fact, when $\mathtt{q}>2$ the phase transition on a $d$-ary tree happens at a certain $\mathtt{p}_{c}(\mathtt{q})$ characterized by being the only (double) root in $(0,1)$ of a certain polynomial $Q_{\mathtt{p},\mathtt{q}}^{(d)}(x)$ (see 2.2). In terms of the recursion (2.5), this correspond to finding $\mathtt{p}$ such that $dg_{\beta,\mathtt{q}}(x)=x$ has only one solution (see Section 2.2.2). For the RCM on a Galton–Watson tree, if $\mathtt{q}>2$, the random recursion (2.5) is still valid but the function $g_{\beta}$ is then not concave anymore, and it is not clear whether a characterization of $p_{c}(\mathtt{q})$ in terms of a certain (averaged?) polynomial remains valid.
@@
The following result is a direct consequence of Theorem 1.8-Proposition 1.9 in the case where $Z\sim\mu$ admits a finite moment of order $s+1$ and of Theorem 1.10 in the case where $\mathbf{E}[Z^{s+1}]=+\infty$. Notice that $\log\varphi_{\mathtt{q}}(x)\sim\mathtt{q}x$ as $x\downarrow 0$, so $\pi_{n}\sim\mathtt{q}^{-1}B_{n}$ when $B_{n}$ goes to 0. Additionally, note that from (2.6) we get that $\kappa_{g}:=\lim_{x\downarrow 0}\frac{1}{x^{2}}(x-g_{\beta}(x))=\frac{2-\mathtt{q}}{2\mathtt{q}}(1-\psi_{\mathtt{q}}(\beta))$ if $\mathtt{q}\in(0,2)$ and $\kappa_{g}:=\lim_{x\downarrow 0}\frac{1}{x^{3}}(x-g_{\beta}(x))=\frac{1}{12}(1-\psi_{\mathtt{q}}(\beta)^{2})$ if $\mathtt{q}=2$.
@@
Let $\beta=\beta_{c}$, with $\beta_{c}$ defined by the relation $m\psi_{\mathtt{q}}(\beta_{c})=1$. We then have the following asymptotic results for $\pi_{n}$. Letting $s=1$ in the case $\mathtt{q}\in(0,2)$ and $s=2$ in the case $\mathtt{q}=2$, we have the following:
  • if $\mathbf{E}[Z^{1+s}]<+\infty$, we have, almost surely and in $L^{1+s}$,
    $\lim_{n\to\infty}n^{1/s}\pi_{n}=\alpha_{\mathtt{q}}W\qquad\text{ with }\ \alpha_{\mathtt{q}}:=\begin{cases}\frac{2}{2-\mathtt{q}}\frac{m}{m-1}\mathbf{E}[W^{2}]^{-1}&\text{ if }\mathtt{q}\in(0,2)\,,\\[3.0pt] \frac{\sqrt{3}}{\sqrt{2}}\frac{m}{\sqrt{m^{2}-1}}\mathbf{E}[W^{3}]^{-1/2}&\text{ if }\mathtt{q}=2\,.\end{cases}$
  • if $c_{1}L(x)x^{1+s-\alpha}\leqslant\mathbf{E}[(Z\wedge x)^{1+s}]\leqslant c_{2}L(x)x^{1+s-\alpha}$ for some $\alpha\in(1,1+s]$, some slowly varying function $L(\cdot)$ and some constants $c_{1},c_{2}>0$, then
@@
Let us also mention that by classical calculations one has $\mathbf{E}[W^{2}]=\frac{1}{m(m-1)}\mathbf{E}[Z(Z-1)]$. Hence, when $\mathtt{q}\in(0,2)$ we can rewrite the constant $\alpha_{\mathtt{q}}$ as $\alpha_{\mathtt{q}}=\frac{2}{2-\mathtt{q}}\frac{m^{2}}{\mathbf{E}[Z(Z-1)]}$, matching that of [2] in the case $\mathtt{q}=1$ of percolation.
@@
In fact, Section 1.3 gives much more precise results (see Theorem 1.8-Proposition 1.9 and Theorem 1.10), that we describe informally as follows: letting $s=1$ if $\mathtt{q}\in(0,2)$ and $s=2$ if $\mathtt{q}=2$, we have
  • if $\mathbf{E}[Z^{1+s}]<+\infty$, we have $\lim_{n\to\infty}(m\psi_{\mathtt{q}}(\beta))^{-n}\pi_{n}=\alpha_{\mathtt{q}}(\beta)W$ for some constant $\alpha_{\mathtt{q}}(\beta)$;
  • if $\mathbf{E}[(Z\wedge x)^{1+s}]=n^{c+o(1)}$ for some constant $c$, then $C\geqslant(m\psi_{\mathtt{q}}(\beta))^{-n}\pi_{n}\geqslant n^{-c^{\prime}}$ as $n\to\infty$}.